OpenStudy (anonymous):

Simplify (1 − sin x)(1 + sin x)

OpenStudy (alekos):

start by expanding the expression

hartnn (hartnn):

or directly using \(\Large (a+b)(a-b) = a^2-b^2\) a=1 b= sin x

OpenStudy (anonymous):

1^2 - sinx^2 or do I factor?

OpenStudy (anonymous):

it would be sin^2x right?

hartnn (hartnn):

no, now you can you the famous trigonometric identity! \(\Large \color{red}{\sin^2 x+\cos^2x =1}\)

hartnn (hartnn):

so what will be \(1-\sin^2 x\) from that ?

OpenStudy (anonymous):

1 - (sin^2)x = (cos^2)x

OpenStudy (anonymous):

I think

hartnn (hartnn):

thats correct! :) \(\Large \cos^2x\) is your final answer :)

OpenStudy (anonymous):

thank you very much

hartnn (hartnn):

welcome ^_^